# Faster Neighborhood Attention: Reducing the $\mathcal{O}(n^{2})$ Cost of Self Attention at the Threadblock Level

### v2

@@ 3.1 Operators
A standard BMM-style attention forward pass (excluding softmax) is comprised of two operations: $QK^{T}$, which produces pre-softmax attention weights ($A$), and $PV$, which applies post-softmax attention weights ($P$) to values ($V$). These operations are different in that the first is a row-major by column-major GEMM, while the other is a row-major by row-major GEMM.
In the case of neighborhood attention, and sliding window attention in general, these will become GEMV problems. In $QK^{T}$, each query token (vector) is multiplied by its neighboring or surrounding key tokens (matrix), and in $PV$, the set of attention weights corresponding to each query token (vector) is multiplied by corresponding value tokens (matrix). We dub the $QK^{T}$ operation “Pointwise-Neighborhood” (PN) and the $PV$ operation “Neighborhood-Neighborhood” (NN), because the same operations can be reused in the backward pass. PN can compute the gradient for post-softmax attention weights ($\nabla P$) when it is fed the output gradient and $V$ instead of $Q$ and $K$. Similarly, NN can compute the gradient for $Q$ ($\nabla Q$), if it is input the pre-softmax attention gradient ($\nabla A$) and $K$ instead of $A$ and $V$. We define a third operator, which can compute gradients for both $K$ and $V$ given the correct input operands: Inverse-Neighborhood (IN). This operation is mostly similar to NN, but differs in the way attention weights are gathered, as well as the number of attention weights. IN may also require loading more attention weights for every token, because unlike in sliding window attention, the relationship between query and context tokens in neighborhood attention is not commutative. In other words, query token at coordinate $i$ attending to context token at coordinate $j$ does not imply that query token at coordinate $j$ attends to context token at coordinate $i$.
BMM-style implementations of standard self attention have a clear edge over neighborhood and sliding window attention implementations, because they are GEMM problems and by extension not as bound by memory as the latter, both of which are GEMV problems. In addition, GEMV problems cannot target specialized matrix multiply-and-accumulate (MMA) cores, such as Tensor Cores. We aim to resolve this issue by formulating all three operators as GEMM problems, in order to better utilize modern hardware accelerators.
@@ 3.2 Batched GEMM NA
Mapping GEMV problems to GEMM problems is straightforward: A subset of individual vectors (a tile at the threadblock level) can be packed into a matrix, and their corresponding individual matrices can be concatenated. Elements in the resulting matrix will have to be scattered according to the original GEMV problem. This approach is not practical unless the sets of column vectors from the individual matrices intersect, which, when eliminated, will reduce the GEMM problem size. In the context of neighborhood and sliding window attention, the aforementioned individual matrices are context tokens, all of which have coordinates in the original token space, and are selected according to the query coordinate. Due to this, as long as query vectors that are packed into a matrix are spatially proximate, their corresponding context tokens will overlap greatly, which reduces the GEMM problem size. We therefore propose modifying a standard batched GEMM kernel as follows in order to implement the above mentioned operators (PN, NN, and IN):
@@
    GEMM tile sizes corresponding to number of tokens (rows and columns) are mapped to multi-dimensional shapes according to the input spatial rank. For example, a GEMM with $64$ rows is mapped to an $8\times 8$ tile for 2-dimensional problems, or to a $4\times 4\times 4$ tile for 3-dimensional problems, but the exact tile shape can either be fixed, or chosen according to problem size.
@@
    Global pointer iterators and accessors are modified to iterate according to the original layout in global memory instead of a standard row or column major matrix layout.
@@
    Attention weights are only scattered to and gathered from global memory, and cannot be accessed as vectors. This especially impacts kernels targeting Tensor Cores, where asynchronous data prefetching becomes impossible.
Figure 4: A simplified illustration of fused neighborhood attention. Queries, and the key-value pair are tiled according to their spatial rank (1-D, 2-D, 3-D), with the latter haloed to include the entire neighborhood for all corresponding queries in the query tile. The haloed KV tile is further optionally tiled into sub-tiles according to the underlying threadblock-scope GEMM shape. Resulting attention weights from the first GEMM are masked according to neighborhood attention parameters, before undergoing online softmax scaling, and going through the second GEMM with the corresponding value sub-tile.
We implemented these concepts by modifying the implicit GEMM kernels in CUTLASS (2.X API) which implement convolution, and successfully implemented 1-dimensional and 2-dimensional kernels that compute the three neighborhood attention operators. Fig. 3 shows an illustration of the GEMM-based PN kernel. 2223-dimensional kernels will be implemented in the future. The first change is relatively inexpensive, but the second change incurs additional indexing logic that may contribute to latency. The final change is a major bottleneck, which we later show cause lower-precision kernels (FP16/BF16) to regress, to the point of providing little to no improvement in latency compared to their full precision (FP32/TF32) counterparts. NN and IN suffer from this issue more significantly, because gathering attention weights by definition does not allow asynchronously prefetching data for one operand (attention weights) in the GEMM mainloop unless the scalar data type meets the minimum access size (32 bits on Ampere). This forces our FP16/BF16 GEMM-based kernels to fall back to standard global reads, which significantly impact achievable latency. To our knowledge, this issue is unavoidable in most cases, and will continue to be a bottleneck as long as attention weights are stored in global memory. Every attention weight corresponds to a query and context coordinate, and even in theory, vectorized access cannot replace the gather/scatter of attention weights unless: 1. window size is known at compile time (which is not very practical especially in 2-D and 3-D problems), and 2. window size is a power of 2 (which it never is in neighborhood attention because window size must be an odd number).
@@ 3.3 Fused NA
We extend our methodology for implementing neighborhood attention operators using batched GEMM kernels to fused attention kernels like Flash Attention [5]. This is motivated by not only the potential to reduce latency and memory footprint, but also to circumvent the bottleneck in the batched GEMM and naive kernels, scattering and gathering attention weights to global memory. Since attention weights are only computed at the threadblock level and never stored in global memory in fused kernels, the bottleneck will simply cease to exist. We started off with FMHA (developed by the xFormers [^3] https://github.com/facebookresearch/xformers/ team), a fused multi-headed attention kernel based on the CUTLASS 2.X API, which can target architectures even older than Ampere (Maxwell, SM50; Volta, SM70; and Turing, SM75.) By carefully applying our methodology for space-aware tiling, neighborhood attention masking, and customized pointer iterators and accessors, we successfully implemented neighborhood attention for 1-D, 2-D, and 3-D problems.  [^4] Only forward pass is implemented in this version. Fig. 4 presents an overview of how our fused kernels function when dealing with multi-axis data.
@@ 3.4 Dilation and causal masking
We implement dilation support through partitioned tensor views. A dilated neighborhood attention problem can be mapped to a set of non-dilated neighborhood attention problems over non-overlapping partitions of the input. All sub-problems can be compute within the same kernel call. We additionally define and implement causal neighborhood attention into our fused kernel, which can be crucial to certain applications where only one spatial dimension requires causal masking (i.e. video embeddings may benefit from causally masked attention across the time axis and standard attention across height and width axes.)
3.5 Auto-tuning
GEMM kernels are, among other settings, parameterized by their tiling shape. Multi-dimensional variants (2-D and 3-D neighborhood attention) can also be parameterized by their fine-grained tile sizes, introduced by our formulation. As mentioned earlier, a GEMM with row tile size 64 can be mapped to a number of valid 2-D and 3-D tiles (i.e. $x\times y$ for all positive integers $x$ and $y$ where $xy=64$, and $x\times y\times z$ for all positive integers $x$, $y$, and $z$ where $xyz=64$.) As a result, selecting tiling sizes based on factors such as problem size, hardware, and environment can further decrease achievable latency. We therefore implement a very simple auto-tuning method as a proof of concept, which is inspired by similar techniques used in packages such as cuDNN. When auto-tuning is activated, a hash map is created and maintained for the lifetime of the application, which maps problems (defined by problem size, data type, and other such factors) to a tiling configuration. Upon encountering problems for the first time, they are benchmarked with a pre-defined set of tiling configurations, which determines and the best configuration, and stores it in the hash map.
@@ 3.6 Limitations
Our formulation of GEMM-based and fused neighborhood attention kernels poses a critical question: how much overhead can one expect from switching from a standard self attention kernel to neighborhood attention? As pointed out in Sec. 3.2, our GEMM-based kernels suffer from a major bottleneck, especially in lower-precision, which stems from the scatter / gather of attention weights, which limits the ability to hide MMA prefetching latency. We consider this to be an unavoidable issue in GEMM-based neighborhood and sliding window attention kernels. Unsurprisingly, our proposed changes to fused implementations are also not free. Changes that we find unavoidable, which in some cases can cause our fused kernels to incur higher latency than the self-attention baseline (FMHA) are the following (ordered by most significant to least significant):
@@
    Kernels specialized for 2-D and 3-D problems fall back to less optimal data iterators which cannot rely on pre-computed pointer offsets as in 1-D problems, and will therefore need to both compute and apply pointer offsets per access. While this is similar to standard practice for data permutation / scattering on the fly in CUTLASS 2.X GEMMs, it is typically less performant than pre-computed offsets. In addition, computing pointer offsets incurs more integer ops with more dimensions, which is repeated per global memory access. We find this to be the most significant contributor to additional latency in our fused kernels, when compared to the baseline fused self attention kernel, FMHA. However, this is usually only significant when window size is approximately equal to input size. Smaller window sizes hide the additional latency by computing fewer attention weights, which reduces both the number of global memory accesses and MMAs.
@@
    The attention masking logic, which depends on corresponding query and context token coordinates, spatial size, and window size, introduces additional integer ops in order to map linear indices to coordinates (unlike in 1-D problems where the mapping is the identity function), which again gets more complicated with more dimensions. This, along with additional statements in the masking condition, contributes to latency, and is expected to worsen with more dimensions. We suspect that it is possible to predicate some of the attention masking logic which would reduce some of the latency, but we leave that for future efforts.
Despite these issues, we find that our fused kernels can still match or outperform our self attention baseline in approximately 97.8% of 1-D, 85.5% of 2-D, and 92.2% of all problem sizes that we benchmarked.
@@ 4 Experiments
Naive - 4.7 % 0.0 % 20.3 % 7.0 %
GEMM 96.4 % - 0.0 % 61.9 % 33.4 %
Fused 100.0 % 100.0 % - 97.8 % 89.8 %
@@
Naive - 34.3 % 8.4 % 38.3 % 17.5 %
GEMM 65.9 % - 0.0 % 54.1 % 23.3 %
Fused 93.7 % 100.0 % - 85.5 % 59.7 %
@@
Naive - - 0.3 % 49.3 % 24.3 %
Fused 99.8 % - - 92.3 % 69.4 %
Table 1: Overview of our FP16 performance benchmark. We benchmark naive neighborhood attention kernels against our proposed GEMM and fused kernels in half precision, over a large set of problem sizes varying in batch size, spatial size, number of attention heads, and dimensions per head, and over different window sizes and dilation values. For every problem size, we also benchmarked self attention running with the xFormers’ FMHA (our baseline) and Flash Attention V2. Note that we have not yet implemented a GEMM kernel for 3-dimensional neighborhood attention.
@@
Naive - 0.1 % 0.0 % 31.2 %
GEMM 99.8 % - 29.9 % 89.4 %
Fused 100.0 % 72.1 % - 94.7 %
@@
Naive - 24.6 % 18.9 % 58.1 %
GEMM 76.0 % - 36.1 % 85.5 %
Fused 81.9 % 66.8 % - 92.9 %
@@
Naive - - 0.2 % 62.9 %
Fused 99.8 % - - 95.2 %
Table 2: Overview of our FP32 performance benchmark. As in Tab. 1, we benchmark naive neighborhood attention kernels against our proposed GEMM and fused kernels in full precision, over a large set of problem sizes varying in batch size, spatial size, number of attention heads, and dimensions per head, and over different window sizes and dilation values. For every problem size, we also benchmarked self attention running with the xFormers’ FMHA (our baseline). Flash Attention V2 is excluded as it does not support full precision. Note that we have not yet implemented a GEMM kernel for 3-dimensional neighborhood attention.
@@
We evaluate the performance of our proposed methods by measuring their latency against existing kernels in $\mathcal{N}ATTEN$. Most use cases in $\mathcal{N}ATTEN$ target naive CUDA kernels, with the exception of 2-D neighborhood attention with 32-dimensional attention heads. $\mathcal{N}ATTEN$ implements tiled kernels for those cases for up to and including window size 13 × 13, and only for the $QK$ operation. However, we treat all kernels in $\mathcal{N}ATTEN$ as our baseline, and will refer to them as naive kernels hereinafter. We run a fixed set of problem sizes that vary in batch size, spatial size, number of attention heads, and dimensions per attention head through every implementation on an NVIDIA A100 GPU and measure their latency. We iterate through multiple neighborhood attention window sizes and dilation values for every problem size. A summary of these benchmarks is presented in Tab. 1 (FP16) and Tab. 2 (FP32). We find that our GEMM-based kernels can improve or match the naive latency in approximately 96% of 1-D problems (of 6150), and 66% of 2-D problems (of 5676) in half precision, and approximately 100% of the 1-D problems and 76% of the 2-D problems in full precision. Note that over 40% of the 2-D problems target tiled kernels in $\mathcal{N}ATTEN$, which we find can sometimes outperform our GEMM-based kernels. Upon re-evaluating our benchmark without the tiled kernels in $\mathcal{N}ATTEN$, our GEMM-based kernels can improve or match 84% of the 2-D problems in half precision and 100% of them in full precision. Another point of disadvantage in our GEMM-based kernels is the memory alignment issue noted in Sec. 3.2, which is especially worsened by both lower-precision and higher spatial rank (i.e. 2-D over 1-D.) On the other hand, our fused kernels improve or match the naive latency in approximately 100% of both 1-D (of 6150) and 3-D problems (of 2448) in both half precision and full precision, while only improving approximately 94% of 2-D problems in half precision and 82% of them in full precision. We again find that if we only compare to non-tiled kernels, fused neighborhood attention kernels improve or match 100% of 2-D problems in both half and full precision. We also find that our fused kernels match or outperform our GEMM kernels in 100% of both 1-D and 2-D problems in half precision, while only doing so in approximately 72% of 1-D problems and 67% of 2-D problems in full precision. Our primary takeaway from this is that despite the mentioned bottleneck in our GEMM kernels, they can still be a useful choice for problems that are heavily compute-bound, which is sometimes the case in full precision. In both Tab. 1 and Tab. 2 we also inspect the percentage of problem sizes in which using our fused neighborhood attention kernel is outperformed by the FMHA kernel. This is only to inspect additional overhead caused by our implementation, which we expect to be more noticeable in 2-D and 3-D problems. Some of the overhead may be avoidable, but our takeaway is that it is unlikely to be fully avoidable, as pointed out in Sec. 3.6.

### v3

@@ 3.1 Operators
[…]
In the case of neighborhood attention, and sliding window attention in general, these will become General Matrix-Vector Multiplication (GEMV) problems. In $QK^{T}$, each query token (vector) is multiplied by its neighboring or surrounding key tokens (matrix), and in $PV$, the set of attention weights corresponding to each query token (vector) is multiplied by corresponding value tokens (matrix). Given that some of these operations can be reused in the backward pass, we dub the $QK^{T}$ operation “Pointwise-Neighborhood” (PN) and the $PV$ operation “Neighborhood-Neighborhood” (NN). PN can compute the gradient for post-softmax attention weights ($\nabla P$) when operating on the output gradient instead of $Q$, and $V$ instead of $K$. Similarly, NN can compute the gradient for $Q$ ($\nabla Q$) when operating on the pre-softmax attention gradient ($\nabla A$) instead of $A$ and $K$ instead of $V$. We define a third operator, which can compute gradients for both $K$ and $V$: Inverse-Neighborhood (IN). This operation is very similar to NN, but differs in gather pattern, as well as the number of attention weights. IN may require loading more attention weights for every token, because unlike in self attention, the relationship between query and context tokens in neighborhood attention is not commutative. In other words, query token at coordinate $i$ attending to context token at coordinate $j$ does not imply that query token at coordinate $j$ attends to context token at coordinate $i$.
BMM-style implementations of standard self attention have a clear edge over neighborhood and sliding window attention implementations, because they are GEMM problems and by extension not as bound by memory bandwidth as the latter, all of which are GEMV problems. In addition, GEMV problems cannot effectively utilize matrix multiply and accumulate (MMA) accelerators, such as Tensor Cores. We aim to minimize this issue by formulating all three operators as batched GEMM problems with scatter/gather fusion, in order to better utilize modern hardware accelerators.
@@ 3.2 Batched GEMM NA
We transform the aforementioned GEMV problems into batched GEMMs with scatter/gather fusion. At an abstract level, implementations of GEMM-based neighborhood attention predicate the execution of tiled MMAs on whether any of the rows in the query tile interact with at least one of the rows in the context tile, given the context window size, dilation, and other masking-related parameters. We propose modifying a CUTLASS GEMM as follows in order to implement PN, NN, and IN:
@@
    GEMM tiling is done according to the original multi-dimensional layout of the token mode in QKV. For example, if the attention problem is 1-D, query and context tensors are tiled along the sequence into tiles of size 64, for a 2-D problem, the token mode, which is comprised of height and width, are tiled by a 2-D tiler of the same size, like 8 × 8.
@@
    Predication logic, and global pointer iterators and accessors are modified to iterate according to the original layout in global memory instead of assuming a standard rank-2 matrix layout.
@@
    Attention weights are required to be scattered to and gathered from global memory, which in 16-bit or lower precision cannot be copied asynchronously (with LDGSTS), which breaks pipelining in those kernels on modern architectures. This is because the minimum transaction size for LDGSTS is 32 bits.
Figure 4: A simplified illustration of fused neighborhood attention. $Q$ and $KV$ tensors are tiled according to their spatial layout (1-D, 2-D, 3-D), with the latter haloed to include the entire neighborhood for all corresponding queries in the query tile. Resulting attention weights from the first GEMM are masked according to neighborhood attention parameters, before undergoing online softmax scaling, and going through the second GEMM with the corresponding value sub-tile.
We implemented these concepts by extending implicit GEMM (convolution) in CUTLASS (2.X API) into kernels that compute the three neighborhood attention operators in 1-D and 2-D. Fig. 3 shows an illustration of the 2-D GEMM-based PN kernel. The first change is relatively inexpensive, but the second change incurs additional predication and indexing logic that can result in additional overhead and register pressure. The final change is a major bottleneck, and leads to lower-precision kernels (FP16/BF16) providing little to no improvement compared to their full precision (FP32/TF32) counterparts. NN and IN suffer from this issue more significantly, because gathering attention weights (LDG) breaks pipelined kernels on Ampere, since they load GEMM operands asynchronously (LDGSTS), which has a minimum transaction size of 32 bits. This forces our FP16/BF16 GEMM-based kernels to fall back to global loads (LDG), which significantly impacts achievable runtime. To our knowledge, this issue is unavoidable in most cases, and will continue to be a bottleneck as long as attention weights are stored in global memory.
@@ 3.3 Fused NA
We extend our methodology for implementing neighborhood attention operators using batched GEMM kernels to fused attention kernels like Flash Attention [6]. This is not only motivated by the potential to reduce runtime and memory footprint, and potentially making neighborhood attention actually bound by compute, but also to circumvent the bottleneck in the batched GEMM and naive kernels: scatter/gathering attention weights to/from global memory. Since attention weights are only computed at the threadblock level and never fully stored in global memory in fused kernels, the bottleneck will simply cease to exist. We started off with xFormers FMHA [13], a fused multi-headed attention kernel based on the CUTLASS 2.X API, which can target architectures even older than Ampere (Maxwell, SM50; Volta, SM70; and Turing, SM75.) By carefully applying our methodology for space-aware tiling, neighborhood attention masking, and software predication for multi-dimensional tensor layouts, we successfully implemented neighborhood attention for 1-D, 2-D, and 3-D problems. Fig. 4 presents an overview of how our fused kernels function when dealing with multi-dimensional (multi-axis) data.
@@ 3.4 Dilation and causal masking
Our methodology allows for dilation support trivially, through simple partitioning and slicing ahead of time. A dilated neighborhood attention problem can be mapped to a set of non-dilated neighborhood attention problems over non-overlapping tiles of the input. All sub-problems can be computed within the same kernel call, simply by issuing more CTAs in the grid. We additionally define and implement causal neighborhood attention into our fused kernel, which can be crucial to certain applications where only one spatial dimension requires causal masking (i.e. video embeddings may benefit from causally masked attention across the time axis and standard attention across height and width axes, which would be an exact 3-D spatio-temporal attention module.)
3.5 Notes on arithmetic intensity
Arithmetic intensity is the ratio of floating point operations over bytes of memory transactions, as defined by the Roofline model [25]:
Arithmetic $\text{Arithmetic Intensity}=\frac{N_{ops}}{N_{bytes}}$ (2)
Arithmetic intensity is typically used to determine whether an implementation/algorithm is bound by memory bandwidth or computational capacity, on a given problem size and hardware. Let’s consider a simplified representation of self attention, where we only look at pure matrix multiplication FLOPs and bytes. Self attention is comprised of two back-to-back BMMs, which would be $2bhn^{2}d$ FLOPs for each of the BMMs, where $b$, $h$, $n$, and $d$ denote batch size, number of attention heads, sequence length, and per-head dimension respectively. In total, that would be $4bhn^{2}d$ FLOPs. In the unfused implementation, 4 tensors with size $bhnd$ ($Q$, $K$, $V$ and output) are accessed in global memory, along with one intermediary tensor with size $bhn^{2}$ (attention weights or $P$), which is accessed twice. In total, that is $(4\times bhnd+2\times bhn^{2})\times s_{dtype}$ bytes, where $s_{dtype}$ is the byte size of the tensor element type. When implemented with fused attention, however, the number of bytes accessed for matrix multiplication from global memory is reduced to only reads and writes for $Q$, $K$, $V$, and attention outputs, or $4\times bhnd\times s_{dtype}$.
Unfused implementations of attention are typically memory-bandwidth-bound at scale, given that their arithmetic intensity approaches a constant value as sequence length grows. If we take the limit of their intensity according to the aforementioned approximation of FLOPs and transaction bytes, as $n\to\infty$ with everything else as constants, we see that:
$\lim_{n\to\infty}\frac{4bhn^{2}d}{(4bhnd+2bhn^{2})s_{dtype}}=\lim_{n\to\infty}\frac{2nd}{(2d+n)s_{dtype}}=\frac{2d}{s_{dtype}}$ (3)
Fused attention therefore solves a key problem here, by reducing the number of memory transactions from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$, which means as sequence length grows, the limit of arithmetic intensity in fused attention does not converge, and it will therefore only become bound by computational capacity. This means that optimal fused attention kernels can almost fully utilize the underlying computational power of modern GPUs, and is the reason behind FP8 attention kernels for the Hopper architecture exceeding the 1 petaFLOP/s threshold [3, 20].
A natural question to ask is what happens to local attention patterns such as neighborhood attention, which promised to deliver more efficiency. In fused implementations of neighborhood attention (i.e. our proposed FNA), we can look at the growth of arithmetic intensity similar to self attention. If we consider the FLOPs for neighborhood attention to be $4bhn\ell d$, where $\ell$ is the size of the attention window, and that the number of global memory transaction bytes is the same as fused self attention (worst case), $4\times bhnd\times s_{dtype}$, then we see that as $n\to\infty$, we converge towards a constant again, therefore making neighborhood attention more memory-bandwidth-bound as $n$ alone scales:
$\lim_{n\to\infty}\frac{4bhn\ell d}{4bhnds_{dtype}}=\frac{\ell}{s_{dtype}}$ (4)
However, the constant here is a function of $\ell$, the size of our attention window, which means that as we scale the sequence length or feature map size, attention window size will determine whether or not the problem is bound by memory bandwidth or computational power. Since smaller window sizes are closer to linear projections, and larger window sizes are closer to self attention, the fact that neighborhood attention can be bound by compute or memory bandwidth depending on window size is not a surprise. Therefore, it is highly recommended to choose neighborhood attention window sizes according to the input size and even hardware to maximize efficiency gain.
@@ 3.6 Limitations
Our formulation of GEMM-based and fused neighborhood attention kernels poses a critical question: how much overhead can one expect from switching from a standard self attention kernel to neighborhood attention? As pointed out in Sec. 3.2, our GEMM-based kernels suffer from a major bottleneck, especially in lower-precision, which stems from scatter/gathering of attention weights. We consider this to be an unavoidable issue in unfused implementations of neighborhood and sliding window attention. Unsurprisingly, our proposed changes to fused implementations are also not free. Changes that we find unavoidable, which in some cases can cause our fused kernels to incur higher runtime than the self attention baseline (xFormers FMHA) are the following (ordered by most significant to least significant):
@@
    Kernels specialized for 2-D and 3-D problems are no longer GEMMs, they are General Tensor-Tensor contractions (GETTs)! Similar to convolution, if the input layout is multi-dimensional, then the GEMM is converted to a special case of GETT. On older GPU architectures, this requires more complicated software predication, which will incur more instructions and heavier register usage, whereas on modern architectures like Hopper, the Tensor Memory Accelerator (TMA) can easily provide hardware predication. Our software predication logic is similar to standard practice for such cases in CUTLASS 2.X GEMMs, and similarly less performant than predication for contiguous matrix layouts. We find this to be the most significant contributor to additional runtime in our fused kernels, when compared to the baseline fused self attention kernel, FMHA. However, FNA is perfectly capable of hiding this additional overhead in many cases, and only falls behind in cases close to self attention (window size is approximately the same as input size.)
@@
    The attention masking logic, which depends on corresponding query and context token coordinates, original layout, and window size, introduces additional indexing logic in order to map linear indices to coordinates (unlike in 1-D problems where the mapping is the identity function), and it gets more complicated with more dimensions. This, along with additional statements in the masking condition, contributes to runtime, and is expected to worsen with more dimensions. Together, these contribute to more serious register spilling than the original 1-D kernel.
Despite these issues, we find that our fused kernels can still match or outperform our self attention baseline in approximately 100% of 1-D, 98.6% of 2-D, and 97.3% of 3-D problem sizes that we benchmarked.
@@ 4 Experiments
Table 1: FP16 forward pass benchmark overview. We benchmark naive neighborhood attention kernels against our proposed GEMM and fused kernels in half precision, over a large set of problem sizes varying in batch size, spatial size, number of attention heads, and dimensions per head, and over different window sizes and dilation values. For every problem size, we also benchmarked self attention running with the xFormers FMHA (our baseline) and Flash Attention V2.
@@
Naive - 1.7 % 0.0 % 21.8 % 8.8 %
GEMM 98.7 % - 0.0 % 72.0 % 54.2 %
Fused 100.0 % 100.0 % - 100.0 % 98.2 %
@@
Naive - 16.4 % 0.0 % 32.9 % 15.8 %
GEMM 84.0 % - 0.0 % 59.3 % 29.8 %
Fused 100.0 % 100.0 % - 98.6 % 92.4 %
@@
Naive - - 0.0 % 43.5 % 20.2 %
Fused 100.0 % - - 97.3 % 87.0 %
Table 2: FP32 forward pass benchmark overview. We benchmark naive neighborhood attention kernels against our proposed GEMM and fused kernels in full precision, over a large set of problem sizes varying in batch size, spatial size, number of attention heads, and dimensions per head, and over different window sizes and dilation values. For every problem size, we also benchmarked self attention running with the xFormers FMHA (our baseline).
@@
Naive - 0.0 % 0.0 % 34.6 %
GEMM 99.9 % - 37.7 % 98.4 %
Fused 100.0 % 64.8 % - 99.9 %
@@
Naive - 11.7 % 5.4 % 52.0 %
GEMM 89.5 % - 28.1 % 92.4 %
Fused 96.0 % 74.0 % - 99.3 %
@@
Naive - - 0.0 % 61.1 %
Fused 100.0 % - - 98.6 %
@@
We evaluate the performance of our proposed methods by measuring their runtime against existing kernels in $\mathcal{N}ATTEN$. Most use cases in $\mathcal{N}ATTEN$ target naive CUDA kernels, with the exception of 2-D neighborhood attention with 32-dimensional attention heads. $\mathcal{N}ATTEN$ implements tiled kernels for those cases for up to and including window size 13 × 13, and only for the $QK$ operation. However, we treat all kernels in $\mathcal{N}ATTEN$ as our baseline, and will refer to them as naive kernels. We use a fixed set of problem sizes that vary in batch size, spatial size, number of attention heads, and dimensions per attention head, and run them through every implementation on an NVIDIA A100 GPU and measure their runtime using CUDA events. We iterate through multiple neighborhood attention window sizes and dilation values for every problem size. A summary of these benchmarks is presented in Tab. 2 (FP16) and Tab. 2 (FP32). We find that our GEMM-based kernels can improve or match the naive runtime in approximately 99% of 1-D problems (of 6150), and 84% of 2-D problems (of 5676) in half precision, and approximately 100% of the 1-D problems and 96% of the 2-D problems in full precision. Note that over 40% of the 2-D problems target tiled kernels in $\mathcal{N}ATTEN$, which we find can sometimes outperform our GEMM-based kernels. Another point of disadvantage in the FP16/BF16 variants of our GEMM-based kernels is using LDGs in pipelined kernels, noted in Sec. 3.2. On the other hand, our fused kernels improve or match the naive runtime in approximately 100% of both 1-D (of 6150) and 3-D problems (of 2448) in both half precision and full precision, an 100% of 2-D problems in half precision, while only improving approximately 96% of 2-D problems in full precision. We also find that our fused kernels match or outperform our GEMM kernels in 100% of both 1-D and 2-D problems in half precision, while only doing so in approximately 65% of 1-D problems and 74% of 2-D problems in full precision, which is not very surprising given that full precision is typically more memory-bandwidth-bound. In both Tab. 2 and Tab. 2 we also inspect the percentage of problem sizes in which using our fused neighborhood attention kernel is outperformed by the FMHA kernel. This is only to inspect additional overhead caused by our implementation, which we expect to be more noticeable in 2-D and 3-D problems. Some of the overhead may be avoidable, but our takeaway is that it is unlikely to be fully avoidable, as pointed out in Sec. 3.6.